# Estimating the strength of Lorentzian distribution in non-commutative geometry by solar system tests

### v3

@@ IV Conclusion and outlook
We conducted detailed calculations of the four classical tests of general relativity in Schwarzschild space-time with a Lorentzian distribution, including planetary orbital precession, light deflection, radar wave delay, and gravitational redshift effects. The study reveals that the non-commutative parameter has a markedly stronger impact on the motion of time-like particles compared to its influence on light rays. With precise experimental observations, the parameter is ultimately constrained within the range $\Theta\leq 0.067579\leavevmode\nobreak\ \mathrm{m}^{2}$, which is given by the observations of Mercury’s precession. This result is consistent with the view that $\sqrt{\Theta}$ is on the order of the Planck length $\ell_{p}$. Moreover, this range exceeds the Planck scale by several orders of magnitude. However, considering that our study focuses on large-scale gravitational experiments within the solar system, it is expected that the derived parameter range will indeed surpass the Planck length. Within the scale of astronomical measurements, this result presents a relatively stringent constraint.
It should be noted that this paper only considers the low-order correction of the non-commutative parameter. Taking into account higher-order corrections of non-commutative geometry would help provide a more accurate understanding of its implications in gravitational theory. Additionally, because the external space-time of the Sun is equivalent to the gravitational field produced by a point mass, the direct application of the Lorentzian distribution modified by point particles is reasonable. However, the Sun itself is a celestial body with a certain volume and a complex mass distribution. If we model the Sun as an perfect fluid, studying the gravitational field generated by an perfect fluid in non-commutative geometry would hold greater practical significance. In addition, studying the Schwarzschild space-time under Gaussian distributions and exploring alternative matter distribution within the framework of non-commutative geometry would also be valuable research topics.

### v4

@@ IV Conclusion and outlook
We conducted detailed calculations of the four classical tests of GR in Schwarzschild space-time with a Lorentzian distribution, including planetary orbital precession, light deflection, radar wave delay, and gravitational redshift effects. The study reveals that the non-commutative parameter has a markedly stronger impact on the motion of time-like particles compared to its influence on light rays. With precise experimental observations, the parameter is ultimately constrained within the range $\Theta\leq 0.067579\leavevmode\nobreak\ \mathrm{m}^{2}$, which is given by the observations of Mercury’s precession. This result is consistent with the view that $\sqrt{\Theta}$ is on the order of the Planck length $\ell_{p}$. Moreover, this range exceeds the Planck scale by several orders of magnitude. However, considering that our study focuses on large-scale gravitational experiments within the solar system, it is expected that the derived parameter range will indeed surpass the Planck length. Within the scale of astronomical measurements, this result presents a relatively stringent constraint.
It should be noted that this paper only considered the low-order correction of the non-commutative parameter. Taking into account higher-order corrections of non-commutative geometry would help provide more accurate understanding of its implications in gravitational theory. Additionally, because the external space-time of the Sun is equivalent to the gravitational field produced by a point mass, the direct application of the Lorentzian distribution modified by point particles is reasonable. However, the Sun itself is a celestial body with a certain volume and a complex mass distribution. If we model the Sun as a perfect fluid, studying the gravitational field generated by a perfect fluid in non-commutative geometry would hold greater practical significance. Besides, when different distributions (such as Gaussian distributions) are considered, the upper limit of the non-commutative parameter $\Theta$ will also change. This paper only discusses the Schwarzschild space-time modified by Lorentz corrections. Moreover, considering the effect of gravitational radiation, the orbits of objects in motion around the gravitational center will continuously shrink. A further discussion of the effects of shrinking orbits will provide more precise constraints on the parameters. However, considering that gravitational effect within the solar system is not significant, this effect is expected to be very weak. Furthermore, it should be noted that in Hyperbolic , researcher has calculated the general relativistic effects on unbound hyperbolic trajectories, where the method used can be generalized to any modified gravity models. Applying this method to non-commutative geometry and further exploring the effects of Lorentzian or Gaussian distributions on hyperbolic trajectories would constitute a novel topic of research.